OpenStudy (anonymous):

What are the center, vertices, foci, and asymptotes of the hyperbola with equation (y 1 1)2 36 2 (x 2 2)2 64 5 1?

OpenStudy (accessdenied):

Welcome to OpenStudy! It appears there was an issue in transferring the problem either by copy paste or by uncertainty of expressing the equation. I am interpreting this as: \( \displaystyle \dfrac{y^2}{36} - \dfrac{x^2}{64} = 1 \) If this is not the equation you were trying to write, please use the equation editor or draw function (both located below the question box) to draw it out! :)